OpenStudy (anonymous):

Consider the curve given by x^2 + sin(xy) + 3y^2 = C, where C is a constant. The point (1, 1) lies on this curve. Use the tangent line approximation to approximate the y-coordinate when x = 1.01. 0.996 1 1.004 Cannot be determined 1.388

OpenStudy (anonymous):

1.004

OpenStudy (dumbcow):

\[f(x) \approx f'(a) (x-a) + f(a)\] a = 1 \[f'(1) = -\frac{2x+y \cos(xy)}{x \cos(xy) +6y} = -\frac{2+\cos(1)}{\cos(1) + 6} \approx -.388407\] \[f(1.01) \approx -.388407 (.01) + 1 = .996\]